OpenStudy (anonymous):

Find the equation of the line perpendicular to the line x-2y=-5 containing the points (4,0)

TheSmartOne (thesmartone):

so -2y=-x-5 y=1/2x+5/2 So the perpendicular line has a slope of -2 y=-2x+b 0=-2(4)+b 0=-8+b b=8 So your final equation is y=-2x+8

OpenStudy (anonymous):

Thanks!